OpenStudy (anonymous):

A polynomial function can be written as (x + 2)(x + 3)(x - 5). What are the x-intercepts of the graph of this function? (2, 0), (3, 0), (-5, 0) (-2, 0), (-3, 0), (5, 0) (2, 0), (3, 0), (5, 0) (-2, 0), (-3, 0), (-5, 0)

OpenStudy (anonymous):

\[x+2=0\iff x=-2\\ x+3=0\iff x=-3\] last one is for you

OpenStudy (anonymous):

let me know what you get

OpenStudy (anonymous):

im lost :(

OpenStudy (anonymous):

really? lets go slow

OpenStudy (anonymous):

i suck at math lol...

OpenStudy (anonymous):

the \(x\) intercept means the value of \(x\) that makes the expression equal to zero, in other words \[(x + 2)(x + 3)(x - 5)=0\] is what you are trying to solve is that much clear?

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

ok then the product of a bunch of numbers can only be zero if one is them is zero so it will be zero if \[x+2=0\]

OpenStudy (anonymous):

So it would be the first one!

OpenStudy (anonymous):

there are three answers , i.e. three numbers that would make t his equal to zero

OpenStudy (anonymous):

if you solve \[x+2=0\] for \(x\), what do you get?

OpenStudy (anonymous):

there could only be one answer lol.. and isn't it (2, 0), (3, 0), (-5, 0)?

OpenStudy (anonymous):

no it isn't when i said there are three answers, i mean there are three numbers what would make this equal to zero i did not mean three of your answer choices

OpenStudy (anonymous):

do you know how to solve \[x+2=0\] for \(x\)? if that is not clear, that is fine, let me know and i will explain

OpenStudy (anonymous):

lol I already submitted my assignment! thank you for helping tho